OpenStudy (fanduekisses):

If the function f(x) is differentiable and f(x)= {ax^3-6x; if x<=1 bx^2+4; x>1 , then a = ?

OpenStudy (anonymous):

take the derivatives plug in 1 set them equal solve for a

jimthompson5910 (jim_thompson5910):

a function f(x) is differentiable if 1) f(x) is continuous AND 2) f ' (x) is continuous

OpenStudy (fanduekisses):

\[f(x) = ax^3-6x;~~if~~x \le1~~~~ bx^2+4;~~~x>1~,then~~a=\]

OpenStudy (anonymous):

although i am a bit confused because there is no a in the second one

OpenStudy (anonymous):

\[f(x)=ax^3-6x\\ f'(x)=?\]

OpenStudy (fanduekisses):

f'(x)= 3ax^2-6

OpenStudy (anonymous):

ok and \[f'(1)=3a-6\]

OpenStudy (anonymous):

\[f(x)=bx^2+4\\ f'(x)=?\]

jimthompson5910 (jim_thompson5910):

Is the function this? \[\Large f(x) = \begin{cases} ax^3-6x \ \ \ \text{if } x \le 1 \\ bx^2+4 \ \ \ \ \ \ \text{if } x>1 \end{cases}\]

OpenStudy (fanduekisses):

f'(x)= 2bx

OpenStudy (anonymous):

right and \[f'(1)=2b\]

OpenStudy (fanduekisses):

yes, how did you do the piecewise like that?

OpenStudy (anonymous):

set \[3a-6=2b\] solve for \(a\) i guess

OpenStudy (anonymous):

if you want to see the code, right click choose "show math as" then "latex"

OpenStudy (freckles):

well you need left limit=right limit of original function also

OpenStudy (freckles):

you will have a system of equations to solve

OpenStudy (anonymous):

ooh right!!

jimthompson5910 (jim_thompson5910):

yeah, freckles is correct, you need to make sure f(x) is continuous at x = 1

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} ax^3-6x& \text{if} & x <1 \\ bx^2+4& \text{if} & x \geq 1 \end{array} \right.\] is another way to make a piecewise function

OpenStudy (anonymous):

plug 1 in to both, set them equal

OpenStudy (anonymous):

take the derivatives of both plug in 1 again and set them equal

OpenStudy (anonymous):

you will have two equations and two unknowns

jimthompson5910 (jim_thompson5910):

satellite73 you have the inequality signs mixed up

OpenStudy (fanduekisses):

the first one is supposed to be if x<=1 and the 2nd x>1

OpenStudy (anonymous):

yeah i know

OpenStudy (anonymous):

the beauty if it makes absolutely no difference

OpenStudy (anonymous):

*is not "if"

OpenStudy (fanduekisses):

ok so I pug in I get: a-6=b+4

OpenStudy (anonymous):

you can almost do it in your head since 1 is so easy to compute with the first equation is \[a-6=b+4\]

OpenStudy (anonymous):

yup

OpenStudy (fanduekisses):

yeah i take the derivative

OpenStudy (anonymous):

ok then plug in 1

OpenStudy (anonymous):

take the derivative of \(ax^2-6a\) as before you did this already

OpenStudy (anonymous):

lol yeah but i made a typo it is \[ax^3-6x\]

OpenStudy (fanduekisses):

3ax^2 - 6 and 2bx

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

plug in 1, set them equal

OpenStudy (fanduekisses):

3a-6=2b

OpenStudy (anonymous):

k almost done

OpenStudy (anonymous):

\[a-6=b+4\\ 3a-6=2b\] solve that system

OpenStudy (fanduekisses):

I got -14

OpenStudy (anonymous):

now it is donkey work solving this i wouldn't give it any attention, i would cheat http://www.wolframalpha.com/input/?i=a-6%3Db%2B4%2C+3a-6%3D2b

OpenStudy (anonymous):

\(a=-14\) is what wolfram got, so you must have done it right (i hope)

OpenStudy (fanduekisses):

yay :) thanks

OpenStudy (anonymous):

yw and don't forget if you ever need code right click on it good for copying and pasting too

OpenStudy (fanduekisses):

how do I type a piece wise function?

jimthompson5910 (jim_thompson5910):

Fanduekisses, right click on this code (shown below) \[\Large f(x) = \begin{cases} ax^3-6x \ \ \ \text{if } x \le 1 \\ bx^2+4 \ \ \ \ \ \ \text{if } x>1 \end{cases}\] go to "show math as" then go to "tex commands", you'll see something like this pop up `\Large f(x) = \begin{cases} ax^3-6x \ \ \ \text{if } x \le 1 \\ bx^2+4 \ \ \ \ \ \ \text{if } x>1 \end{cases}` you would type that out and surround it with `\[ \]` to have the server process your code and turn it into LaTex. You can also use the equation button below the text box